OpenStudy (anonymous):

A middle school when what watching. The cost was $8 per student and $10 adult. The total cost for the whale watching was $518 for 63 tickets. How many student went whale watching? How many adults?

OpenStudy (anonymous):

let x = number of students, y = number of adults ; set up two equations using the information you have, that is: 8x+10y = 518 and x+y = 63 , now solve for x or y and plug in

OpenStudy (anonymous):

The above poster is right in using a system of equations, but I will show you how to do it with a single variable. Let x be the number of adults and 63-x be the number of students, since those who are not adults must be students. So, 10x+8(63-x)=518. Simplify this to 10x+504-8x=518, simplify to 2x=14, x=7---so 7 adults went and 56 students went :).

OpenStudy (anonymous):

thank you ryclark and Sujay, very helpful!